OpenStudy (anonymous):

Check my answers please: Given f(x) = x²+2 ; g(x) = 3x a)g*f(x): >My answer: g(x²+2) = 3(x²+2) b)f*g(x): >My answer: (3x)²+2 = 3x²+2 c)g^-1*g(x) >My answer: 3x/3 **I'm unsure of B, C.

OpenStudy (anonymous):

b is 9x+2

OpenStudy (anonymous):

B) f*g(x) = 9x + 2 C) y=3x x=3y so, y=x/3 therefore, 3(x/3) = x g^-1*g(x) = x

OpenStudy (anonymous):

@MathMind In c you made y = 3x, the reason that i was x/3 is because isn't that the inverse of 3x=y?

OpenStudy (anonymous):

a) is correct b) is incorrect c) checking...

OpenStudy (anonymous):

c is correct

OpenStudy (anonymous):

b should be (3x)^2+2=9x^2+2

OpenStudy (anonymous):

well yea, i didn't see you didn't simplified it, sorry. but it's right

OpenStudy (anonymous):

and Fruitless is right about b, i miss the (^2) on the x B) \[9x^2+2\]

OpenStudy (anonymous):

cool. Thanks Guys/Girls. :) *Oh and if possible can one of you answer my Inverse question below?*